# Physics-Informed GNN for Medium-High Voltage AC Power Flow with Edge-Aware Attention and Line Search Correction Operator

### v1

@@ Abstract
Index Terms—  AC power flow, physics-informed neural networks, graph neural networks, backtracking line search, power systems
@@ 1 Introduction
State of the practice PF solvers only partly meet these needs. Newton–Raphson (NR) often converges rapidly when well initialized, but calculating inverse of Jacobian via sparse triangular solves is not naturally GPU-friendly [4]. Fast-decoupled load flow (FDLF) is quicker but relies on reactance dominant, weak-coupling assumptions and loses accuracy when those assumptions break [5]. In both approaches, performance hinges on large sparse linear systems where cost worsen with network size and conditioning[6], creating a brittle accuracy–speed trade-off: improving the robustness of the solver often slows it down, whereas making it fast typically degrades accuracy.
@@
Building on this foundation, the pursuit of accuracy at parity speed remains a central research direction, with several proposed architectures and training strategies emerging in the literature. On the architectural side, many GNN variants, especially attention-based methods[19, 20], have been explored; however, two limitations persist: (i) edge physics are encoded only implicitly through node embeddings, and (ii) training relies on NR labels. To address these issues, we propose an edge-aware attention mechanism tailored to a physics-informed architecture; each head is biased by admittance features, producing sharper, physics-faithful message weights.
Beyond architectural design, another limitation of existing PIGNNs is that the physics loss guiding training remains inoperative during inference, leaving the forward unroll without a built-in convergence guarantee on out-of-distribution cases. To provide inference-time globalization, we incorporate a backtracking line search  [21] on a merit function to adaptively select the step size for the $K$ message-passing updates in $(\Delta V,\Delta\theta)$, guaranteeing sufficient decrease of the merit at each iteration.
@@ 2.1 Problem Formulation
J{J} depends nonlinearly on admittances, voltages, and angle differences; building and solving it is expensive on large meshed grids.[25]
@@
$\Delta x\approx\mathcal{G}_{\theta}\!\big(x,Y)\,r(x),$
@@ 2.3 Physics-Informed, Edge-Aware Multi-Head Self-Attention
The edge-dependent bias $\beta_{ij}^{(h)}$ injects admittance features directly into the attention scores, yielding a strong line-physics inductive bias; by contrast, MLP or standard graph attention network [28] aggregation incorporates admittance only implicitly via node features $\Delta P,\Delta Q$.
@@ 4.1 Data Preprocessing and Implementation Details
PIGNN-MLP uses 4 aggregation channels and an update MLP with hidden size 16. PIGNN-Attn uses the same hidden size with 4 attention heads and 1 attention layer per correction step; the remaining settings are identical. Line search uses initial step $\alpha_{0}=1.0$, Armijo constant $c_{1}=10^{-4}$, shrink factor $\beta=0.5$, and minimum step $\alpha_{\min}=5\times 10^{-2}$. Per-iteration caps are $\Delta\theta_{\max}=0.3$ rad and $\Delta|V|_{\max}=10\%$ of the current magnitude, with voltage bounds $V_{\min}=0.8$ and $V_{\max}=1.2$. Training is unsupervised with discounted physics loss $\gamma=0.9$, $K=40$ message-passing steps, AdamW with weight decay $10^{-3}$, cosine learning-rate annealing from $10^{-4}$ to $10^{-6}$ every 20 epochs, and batch size 64 with block-diagonal batching; runs use a single RTX 3070 8 GB GPU and an Intel Core i7–12700 with 12 cores and 20 threads, up to 4.9 GHz. Hyperparameters were selected by a compact grid search that varied aggregation channels at 4, 8, 16, 32; hidden size from 16 to 128; message-passing steps from 10 to 80 in steps of 10; attention heads 4 or 8; attention layers 1 to 4; angle caps 0.3 to 3.14 rad; magnitude caps 10% to 100% of $V$; and two line search variants: steeper decrease with $\alpha_{0}=1.0$, $c_{1}=10^{-3}$, $\beta=0.5$, $\alpha_{\min}=5\times 10^{-2}$; conservative monotone with $\alpha_{0}=0.5$, $c_{1}=10^{-4}$, $\beta=0.7$, $\alpha_{\min}=5\times 10^{-2}$. The reported configuration is best by validation physics loss; wall-clock time breaks ties.
@@ 4.2 Ablation Studies
For HV+MV, using smaller batches improves accuracy, by reducing cross-regime interference from mixed block-diagonal batching. Although trained on $N\!=\!4\text{--}32$ grids, the model scales effectively to much larger HV grids (e.g., up to $N=1024$) with comparable RMSE (about $V\!\approx\!5{\times}10^{-4}$, $\theta\!\approx\!10^{\circ}$).

### v2

@@ Abstract
Index Terms—  AC power flow, physics-informed neural networks, known operator learning, graph neural networks, backtracking line search
@@ 1 Introduction
State of the practice PF solvers only partly meet these needs. Newton–Raphson (NR) often converges rapidly when well initialized, but calculating inverse of Jacobian via sparse triangular solves is not naturally GPU-friendly [25]. Fast-decoupled load flow (FDLF) is quicker but relies on reactance dominant, weak-coupling assumptions and loses accuracy when those assumptions break [22]. In both approaches, performance hinges on large sparse linear systems where cost worsens with network size and conditioning[24], creating a brittle accuracy–speed trade-off: improving the robustness of the solver often slows it down, whereas making it fast typically degrades accuracy.
@@
Building on this foundation, the pursuit of accuracy at parity speed remains a central research direction. A recurring gap is that physics-informed often enters only through the training objective: as a regularizer, it penalizes violations of the physical equations during training, but it does not provide a mechanism that guarantees physically plausible behavior at inference time [18, 14]. Motivated by known-operator learning [18], we inject admittance-derived coupling into message passing via attention biases, yielding an explicit, differentiable known-operator layer inside the computational graph rather than treating admittances as additional input features.
To further operationalize this bias during inference, we augment the unrolled solver with a backtracking line search [1] on a power-mismatch merit function. It adaptively selects the step size $\alpha$ for each of the $K$ updates in $(\Delta V,\Delta\theta)$, enforcing sufficient decrease and improving robustness under out-of-distribution operating conditions.
@@ 2.1 Problem Formulation
J{J} depends nonlinearly on admittances, voltages, and angle differences; building and solving it is expensive on large meshed grids[4].
@@
$\Delta x\approx\mathcal{G}_{\theta}\!\big(x,Y)\,r(x).$
@@ 2.3 Physics-Informed, Edge-Aware Multi-Head Self-Attention
The edge-dependent bias $\beta_{ij}^{(h)}$ injects admittance features into attention scores. From a known-operator learning viewpoint [18], it acts as a lightweight, fully differentiable known-operator layer that makes edge coupling explicit in the forward pass, unlike MLP/GAT-style aggregation [28], where it is only implicit via node features.
@@ 4.1 Data Preprocessing and Implementation Details
PIGNN-MLP uses 4 aggregation channels and an update MLP with hidden size 16. PIGNN-Attn uses the same hidden size with 4 attention heads and 1 attention layer per correction step; the remaining settings are identical. Line search uses initial step $\alpha_{0}=1.0$, Armijo constant $c_{1}=10^{-4}$, shrink factor $\rho=0.5$, and minimum step $\alpha_{\min}=5\times 10^{-2}$. Per-iteration caps are $\Delta\theta_{\max}=0.3$ rad and $\Delta|V|_{\max}=10\%$ of the current magnitude, with voltage bounds $V_{\min}=0.8$ and $V_{\max}=1.2$. Training is unsupervised with discounted physics loss $\gamma=0.9$, $K=40$ message-passing steps, AdamW with weight decay $10^{-3}$, cosine learning-rate annealing from $10^{-4}$ to $10^{-6}$ every 20 epochs, and batch size 64 with block-diagonal batching; runs use a single RTX 3070 8 GB GPU and an Intel Core i7–12700 with 12 cores and 20 threads, up to 4.9 GHz. Hyperparameters were selected by a compact grid search that varied aggregation channels at 4, 8, 16, 32; hidden size from 16 to 128; message-passing steps from 10 to 80 in steps of 10; attention heads 4 or 8; attention layers 1 to 4; angle caps 0.3 to 3.14 rad; magnitude caps 10% to 100% of $V$; and two line search variants: steeper decrease with $\alpha_{0}=1.0$, $c_{1}=10^{-3}$, $\rho=0.5$, $\alpha_{\min}=5\times 10^{-2}$; conservative monotone with $\alpha_{0}=0.5$, $c_{1}=10^{-4}$, $\rho=0.7$, $\alpha_{\min}=5\times 10^{-2}$. The reported configuration is best by validation physics loss; wall-clock time breaks ties.
@@ 4.2 Ablation Studies
For HV+MV, using smaller batches improves accuracy, by reducing cross-regime interference from mixed block-diagonal batching. Although trained on $N\!=\!4\text{--}32$ grids, the model scales effectively to much larger dense HV grids (e.g., up to $N=1024$) with comparable RMSE (about $V\!\approx\!5{\times}10^{-4}$, $\theta\!\approx\!10^{\circ}$).